OpenStudy (anonymous):

Find the difference. 4d2e - d2e A. 3 B. 3de C. 3d2e D. 4d2e

OpenStudy (mathstudent55):

keep in mind that -d^2e = -1d^2 just like x = 1x

OpenStudy (mathstudent55):

For example, 5x^2 - x^2 = 5x^2 - 1x^2 = 4x^2

OpenStudy (anonymous):

i have choices to pick from im not really good at math at all

OpenStudy (mathstudent55):

Think of d^2e as being orenges. What is 42 oranges minus an orange?

OpenStudy (anonymous):

minus 1 orange only that would be 41 .

OpenStudy (mathstudent55):

Right. Since minus an orange means minus 1 orange, the answer is 42 oranges - 1 orange = 41 oranges

OpenStudy (anonymous):

i still do not understandd ...

OpenStudy (mathstudent55):

In this problem, instead of oranges, you have d^2e. d is a variable and e is another variable. You have 42 d^2e - 1 d^2e that leaves 41 d^2e

OpenStudy (mathstudent55):

Sorry. I was looking at the problem and thought it was 42, not 4. I thought the 2 followed the 4, instead of following the d. 4d^2e - d^2e = 4d^2e - 1d^2 = 3d^2e

OpenStudy (mathstudent55):

Use the oranges example with 4 oranges minus 1 orange equals 3 oranges.

OpenStudy (mathstudent55):

This is what it should look like using exponents: